OpenStudy (anonymous):

pls prove that-- sin(A+B)sin(A-B)+sin(B+C)sin(B-C)+sin(C+A)-sin(C-A)=0

OpenStudy (anonymous):

myininaya will do this

OpenStudy (anonymous):

sin(a+b)sin(a-b): \[(\sin(a)\cos(b)+\sin(b)\cos(a))(\sin(a)\cos(b)-\sin(b)\cos(a)) \] \[= \sin^2(a)\cos^2(b)-\sin^2(b)\cos^2(a)\] each one of the terms will become a difference of two squares. i dont know if this leads to a solution or not, seems like a decent start though...gotta read >.<

OpenStudy (anonymous):

thank u joe let me go once more with it

OpenStudy (anonymous):

will u pls say the steps

OpenStudy (anonymous):

ok thank u myininaya

myininaya (myininaya):

thats blank

myininaya (myininaya):

lol

OpenStudy (anonymous):

yeah........!!!!!!

OpenStudy (anonymous):

ha...ha...

OpenStudy (anonymous):

thanks a lot myininaya i ll go through it

myininaya (myininaya):

hey i made a mistake in that attachment

OpenStudy (anonymous):

bt i could not reach the answer i ll post another qstion pls anwser

myininaya (myininaya):

i;m still working on it

myininaya (myininaya):

so i have this so far \[\cos^2B(\sin^2A-\sin^2C)-\sin^2B(\cos^2A-\cos^2C)+2cosCsinA\]

myininaya (myininaya):

so this is s trig class?

OpenStudy (anonymous):

yeah.........

OpenStudy (anonymous):

i hav a doubt that whether i had any mistake in writing the qstion ie, instead of sin(A+B)sin(A-B)+sin(B+C)sin(B-C)+sin(C+A)-sin(C-A)=0, it is sin(A+B)sin(A-B)+sin(B+C)sin(B-C)+sin(C+A)sin(C-A)=0

myininaya (myininaya):

oh no really lol

myininaya (myininaya):

well then i got it lol

OpenStudy (anonymous):

just funny !!!!

myininaya (myininaya):

the parts where i X out just pretend that part is sin(c+a)sin(c-a)

OpenStudy (anonymous):

ya i got it......... thank u for ur concern

myininaya (myininaya):

the trick was to expand all of them and multiply then i wrote everything in terms of sin by using identities mainly sin^2x+cos^2x=1 identity.